# Boundary-Mediated Phases of Self-Propelled Kuramoto Particles

### 13 Mar 2026 v1

@@ II Numerical methods
where $r_{ij}$ is the distance between particle $i$ and particle $j$ and $\sigma_{ij}=(\sigma_{i}+\sigma_{j})/2$ is the average particle diameter. The other force terms in the equation describe the active force, $f_{0}\hat{n}_{i}$, and the interaction with the boundary, $\vec{F}^{w}_{i}$. We set the energy scale of the potential to $\varepsilon=1$ and define $\sigma\sqrt{m/\varepsilon}$ as the unit of time, $\tau_{i}$. The drag coefficient is set to $\gamma=10^{2}$ such that the inertial timescale $m/\gamma$ is negligible. The magnitude of the active force, $f_{0}$, is fixed to $2\varepsilon/\sigma$ well below the maximum steric repulsion ($24\varepsilon/\sigma$ when $r_{ij}=\sigma_{ij}$), ensuring that collision events are dominated by excluded-volume interactions. The orientation of the active force acting on particle $i$ is defined by a unit vector $\hat{n}_{i}(t)=(\cos\theta_{i}(t),\sin\theta_{i}(t))$, where the angle $\theta_{i}(t)$ evolves following the equation
@@
Here, the first term on the r.h.s. of the equation is a scalar white noise, where $\tau_{p}$ is the persistence time, i.e. the typical time needed to change the direction of self-propulsion, and $\xi_{i}(t)$ is a Gaussian random process of mean zero and correlation $\langle\xi_{i}(t)\xi_{i}(t^{\prime})\rangle=\delta(t-t^{\prime})$. The second term is the Kuramoto interaction, where the timescale $\tau_{K}=\pi R_{K}^{2}/K$ is the alignment time, with $K$ being the coupling strength and $R_{K}$ the interaction radius. In the following, we will vary $\tau_{K}$ by varying the coupling strength $K$, while keeping the interaction radius equal to $1.5\sigma$. Particles align by minimizing the sine of the difference between their velocity orientations, namely $\varphi_{v,i}$, which is initialized by setting random velocities of magnitude $v_{0}=f_{0}/\gamma$. In the absence of the Kuramoto term, the model reduces to a system of Active Brownian particles which may acquire local velocity alignment from persistent collisions in the MIPS regime [16, 7]. However, this phenomenon is observed in much denser systems than those studied here, so we can attribute alignment solely to the Kuramoto term. Eqs. (1) and (3) are numerically integrated using the Euler-Maruyama scheme with a time step $\Delta t=5\times 10^{-5}$.
@@
Figure 1: (a) A particle before and after collision with a smooth boundary. (b) Particle-monomer interaction in a rough boundary with radial $f_{r}$ and tangential $f_{\theta}$ components.
@@ III Phases of confined Self-Propelled Kuramoto Particles
By varying $\tau_{p}$ and $\tau_{K}$, we identify three distinct phases reported in Fig. 2. At short persistence time and weak alignment (large $\tau_{K}$), particles reorient more rapidly than they can align with their neighbors, leading to a homogeneous, weakly correlated gas of self-propelled particles. As $\tau_{p}$ increases at fixed $\tau_{K}$, particles are pushed to travel longer along a fixed direction before reorienting. As a consequence, they keep hitting the boundary and spend more time near it. This increases the chances of aligning with their neighbors if the strength of the Kuramoto term is comparable to the magnitude of self-propulsion. Notably, for sufficiently small $\tau_{K}$, or sufficiently strong alignment, a delocalized clustered (DC) phase emerges where particles form a single annular cluster (roughly $2\sigma$ thick) that spans the whole circumference of the wall and rotates around its center. As alignment increases ($\tau_{K}$ decreases) while keeping $\tau_{p}$ fixed, particles are forced to sit at the interaction distance $R_{K}$ leading to the localized clustered (LC) phase where particles form one or more compact aggregates that slide along the boundary. Delocalized and localized clusters select a direction of rotation, either clockwise or counterclockwise, thus spontaneously breaking the symmetry of the confining geometry. We note that, when persistence is low ($\tau_{p}<1$), the cluster direction of motion rapidly reorients and clusters are not bound to slide along the surface, see Video 3 in [1]. We verified the occurrence of the different phases in a larger system composed of $N=32768$ particles and observed no significant variations.
@@
This direct link between $\tilde{I}$ and the cluster radius provides an estimate of the area fraction, or density, within the cluster. For example, the configuration reported in Fig. 2c corresponds to $\tilde{I}\approx 0.74$, and by inverting (6) we get $R_{c}/R\approx 0.15$, i.e. a cluster radius of about $15\%$ of the wall radius. Thus, computing the number density as $n_{c}=N/\pi R_{c}^{2}$ yields an area fraction equal to $n_{c}a_{p}\approx 0.61$, where $a_{p}$ is the average particle area. Given the ratio between kinetic and potential energy scales, $f_{0}\sigma/\varepsilon=2$, this density suggests that clusters are in a liquid-like state.
@@ V Conclusions
Our study proves that boundary friction mediates phases of self-propelled aligning objects. This indicates that the underlying mechanism giving rise to collective motion in confined active systems can be inferred from the macroscopic structures and dynamical patterns that emerge near mechanical obstacles. Delocalized structures signal self-propulsion-dominated motility as the primary driving force, exemplified by E. coli bacteria moving via run-and-tumble motion or polar Janus particles [6], and Quincke rollers [35]. By contrast, compact structures underscore the presence of strong neighbor-to-neighbor coordination, as expected in alignment–regulated systems, e.g. bird flocks, swarms, fish schools. In addition, we show that boundary friction can act as a regulatory switch, enabling the system to adapt its accumulation patterns in response to external modifications, such as surface coatings in microfluidic essays of active nematics [21], and rough geometries in experiments of granular active matter [14]. Our results establish a framework for future studies on active systems that are mechanically coupled to their environment, where accumulation near boundaries may trigger rotation as for bacteria in ratchet geometries [15] or deformation as in neural crest cells during development [29].

### 16 Mar 2026 v2

@@ II Numerical methods
where $r_{ij}$ is the distance between particle $i$ and particle $j$ and $\sigma_{ij}=(\sigma_{i}+\sigma_{j})/2$ is the average particle diameter. The other terms in the equation describe the active driving, $v_{0}\hat{n}_{i}$, and the interaction with the boundary, $\mu\vec{F}^{w}_{i}$. We set the energy scale of the potential to $\varepsilon=1$ and define $\sigma\sqrt{m/\varepsilon}$ as the unit of time, $\tau_{i}$. The mobility is set to $\mu=10^{-2}\tau_{i}/m$ such that the inertial relaxation time $m\mu$ is negligible. The magnitude of the active velocity $v_{0}$ is fixed to $2\sigma/\tau_{i}$ corresponding to a kinetic energy scale $mv_{0}^{2}/2=2\varepsilon$, comparable to the interaction energy scale $\varepsilon$. Particle velocity components are initialized from a Gaussian distribution with zero mean and variance $v_{0}^{2}/2$. The orientation of the active force acting on particle $i$ is defined by a unit vector $\hat{n}_{i}(t)=(\cos\theta_{i}(t),\sin\theta_{i}(t))$, where the angle $\theta_{i}(t)$ evolves following the equation
@@
Here, the first term on the r.h.s. of the equation is a scalar white noise, where $\tau_{p}$ is the persistence time, i.e. the typical time needed to change the direction of self-propulsion, and $\xi_{i}(t)$ is a Gaussian random process of mean zero and correlation $\langle\xi_{i}(t)\xi_{i}(t^{\prime})\rangle=\delta(t-t^{\prime})$. The second term is the Kuramoto interaction, where the timescale $\tau_{K}=\pi R_{K}^{2}/K$ is the alignment time, with $K$ being the coupling strength and $R_{K}$ the interaction radius. In the following, we will vary $\tau_{K}$ by varying the coupling strength $K$, while keeping the interaction radius equal to $1.5\sigma$. Particles align by minimizing the sine of the difference between their velocity orientations, namely $\varphi_{v,i}=\arctan(v_{i,y}/v_{i,x})$. In the absence of the Kuramoto term, the model reduces to a system of Active Brownian particles which may acquire local velocity alignment from persistent collisions in the MIPS regime [16, 7]. However, this phenomenon is observed in much denser systems than those studied here, so we can attribute alignment solely to the Kuramoto term. Eqs. (1) and (3) are numerically integrated using the Euler-Maruyama scheme with a time step $\Delta t=5\times 10^{-5}$.
@@
Figure 1: (a) A particle before and after collision with a smooth boundary. (b) Particle-monomer interaction in a rough boundary with radial $f_{r}$ and tangential $f_{\phi}$ components.
@@ III Phases of confined Self-Propelled Kuramoto Particles
By varying $\tau_{p}$ and $\tau_{K}$, we identify three distinct phases reported in Fig. 2. At short persistence time and weak alignment (large $\tau_{K}$), particles reorient more rapidly than they can align with their neighbors, leading to a homogeneous, weakly correlated gas of self-propelled particles. As $\tau_{p}$ increases at fixed $\tau_{K}$, particles travel longer along a fixed direction before reorienting. This enhances collisions between neighboring particles, promoting alignment and cluster formation. Once formed, clusters of aligned particles may reach the boundary and remain localized there due to their net persistent motion. Notably, for sufficiently small $\tau_{K}$, or sufficiently strong alignment, a delocalized clustered (DC) phase emerges where particles form a single annular cluster (roughly $2\sigma$ thick) that spans the whole circumference of the wall and rotates around its center. As alignment increases ($\tau_{K}$ decreases) while keeping $\tau_{p}$ fixed, particles are forced to sit at the interaction distance $R_{K}$ leading to the localized clustered (LC) phase where particles form one or more compact aggregates that slide along the boundary. Delocalized and localized clusters select a direction of rotation, either clockwise or counterclockwise, thus spontaneously breaking the symmetry of the confining geometry. We note that, when persistence is low ($\tau_{p}<1$), the cluster direction of motion rapidly reorients and clusters are not bound to slide along the surface, see Video 3 in [1]. We verified the occurrence of the different phases in a larger system composed of $N=32768$ particles and observed no significant variations.
@@
This direct link between $\tilde{I}$ and the cluster radius provides an estimate of the area fraction, or density, within the cluster. For example, the configuration reported in Fig. 2c corresponds to $\tilde{I}\approx 0.74$, and by inverting (6) we get $R_{c}/R\approx 0.15$, i.e. a cluster radius of about $15\%$ of the wall radius. Thus, computing the number density as $n_{c}=N/\pi R_{c}^{2}$ yields an area fraction equal to $n_{c}a_{p}\approx 0.61$, where $a_{p}$ is the average particle area. Given the ratio between kinetic and potential energy scales, $mv_{0}^{2}/2\varepsilon=2$, this density suggests that clusters are in a liquid-like state.
@@ V Conclusions
Our study proves that boundary friction mediates phases of self-propelled aligning objects. This indicates that the underlying mechanism giving rise to collective motion in confined active systems can be inferred from the macroscopic structures and dynamical patterns that emerge near mechanical obstacles. Delocalized structures signal self-propulsion-dominated motility with weak inter-particle coordination, as observed in E. coli performing run-and-tumble motion and self-propelled Janus particles in the dilute regime [6]. By contrast, when self-propulsion is couple with neighbor-to-neighbor alignment, cohesive collective structures emerge—a mechanism whose universality transcends length scales, from synthetic Quincke rollers [35] to large animal groups, e.g. bird flocks, insect swarms and fish schools. In addition, we show that boundary friction can act as a regulatory switch, enabling the system to adapt its accumulation patterns in response to external modifications, such as surface coatings in microfluidic essays of active nematics [21], and rough geometries in experiments of granular active matter [14]. Our results establish a framework for future studies on active systems that are mechanically coupled to their environment, where accumulation near boundaries may trigger rotation as for bacteria in ratchet geometries [15] or deformation as in neural crest cells during development [29].